# Coloring outside the lines: Spectral bounds for generalized hypergraph colorings

### v1

@@ 6 Edge-coloring hypergraphs
In this final section we shall shine a spectral light on edge-colorings of hypergraphs (Definition 2.26). Here, edges rather than vertices are assigned colors, under the constraint that intersecting edges must receive distinct colors.
@@
Recall that $L=D^{-1}\mathcal{I}\mathcal{I}^{\top}$. Since matrices of the form $AB$ and $BA$ have the same non-zero spectrum, this applies in particular to $L$ and $L^{1}$. Hence, as shown in if we let $m_{V}$ and $m_{E}$ be the multiplicity of the eigenvalue 00 of $L$ and $L^{1}$, respectively, then
@@
[…]
=$\overset{(i)}{=}\frac{N}{M}+(\chi^{\prime}-1)\biggl{(}\frac{N}{M}-\textrm{RQ}(\gamma_{12})\biggr{)}$
@@
Let $\Gamma$ be a $c$-uniform hypergraph in which all vertices are inputs in all edges they belong to, and for which the bound from (24) is attained with equality. Then, we have the following two cases:
@@
    If $\mu_{1}=0$, then $\Gamma$ is $k$-regular for some $k$, and $\chi^{\prime}=k$. Furthermore, the multiplicity of the eigenvalue 00 is at least $k-1$.
@@
Furthermore, if $\mu_{1}=0$, then the functions $\gamma_{1i}$ for $i=1,\ldots,\chi^{\prime}$ from (25) are linearly independent eigenfunctions with eigenvalue $\mu_{1}=0$. Therefore, the multiplicity of 00 is at least $\chi^{\prime}-1=k-1$.
@@
Let $\Gamma$ be a $c$-uniform hypergraph in which all vertices are inputs in all edges they belong to. If $M>N$, then $\mu_{1}=0$, and the bound from (24) becomes $\chi^{\prime}\geq\overline{\deg}$, which does not provide any additional information beyond the trivial bound $\chi^{\prime}\geq\max\deg$. However, if $M<N+\chi^{\prime}-1$, then 00 is an eigenvalue of $L$ with multiplicity at least $N+\chi^{\prime}-1-M$.

### v2

@@ 6 Edge-coloring hypergraphs
In this final section, we shall shine a spectral light on edge-colorings of hypergraphs (Definition 2.26). Here, edges rather than vertices are assigned colors, under the constraint that intersecting edges must receive distinct colors.
@@
Recall that $L=D^{-1}\mathcal{I}\mathcal{I}^{\top}$. Since matrices of the form $AB$ and $BA$ have the same non-zero spectrum, this applies in particular to $L$ and $L^{1}$. Hence, as shown in [21, Corollary 14], if we let $m_{V}$ and $m_{E}$ be the multiplicity of the eigenvalue 0 of $L$ and $L^{1}$, respectively, then
@@
$\overset{\text{WLOG}}{\leq}\frac{N}{M}+2(\chi^{\prime}-1)\frac{\sum_{e\in V_{1}}\sum_{e^{\prime}\in V_{2}}\biggl(\sum_{\begin{subarray}{c}v\in V\\ v\in e\cap e^{\prime}\end{subarray}}\frac{1}{\deg v}\biggr)}{\sum_{e\in V_{1}\cup V_{2}}1}$
$\overset{(i)}{=}\frac{N}{M}+(\chi^{\prime}-1)\biggl(\frac{N}{M}-\textrm{RQ}(\gamma_{12})\biggr)$
@@
Let $\Gamma$ be a $c$-uniform hypergraph in which all vertices are inputs in all edges they belong to, and for which the bound in (25) is attained with equality. Then, we have the following two cases:
@@
    If $\mu_{1}=0$, then $\Gamma$ is $k$-regular for some $k$, and $\chi^{\prime}=k$. Furthermore, the multiplicity of the eigenvalue 0 is at least k1k-1.
@@
Furthermore, if $\mu_{1}=0$, then the functions $\gamma_{1i}$ for $i=1,\ldots,\chi^{\prime}$ from (26) are linearly independent eigenfunctions with eigenvalue $\mu_{1}=0$. Therefore, the multiplicity of 0 is at least $\chi^{\prime}-1=k-1$.
@@
Let $\Gamma$ be a $c$-uniform hypergraph in which all vertices are inputs in all edges they belong to. If $M>N$, then $\mu_{1}=0$, and the bound from (25) becomes $\chi^{\prime}\geq\overline{\deg}$, which does not provide any additional information beyond the trivial bound $\chi^{\prime}\geq\max\deg$. However, if $M<N+\chi^{\prime}-1$, then 0 is an eigenvalue of $L$ with multiplicity at least $N+\chi^{\prime}-1-M$.